answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

An expression of possible loss adverse outcome or negative consequence in terms of probability and severity?

Which term most closely matches the desription: Any real or potential condition that can cause injury, illness, or death to personnel; damage to or loss of equipment or property; degradation of mission capability or impact to mission accomplishment; or damage to the environment.

What is event of Alinghi?

Alinghi is a team formed in 2000 that competes in the America's Cup yacht race. The team was founded by Ernesto Bertarelli, and it won the competition in 2003.

What is the probability that a randomly chosen letter is a vowel?

5/24, or five out of twenty four, or with numbers, five out of infinity

Why estimating before you divide 624 divided by 6 helps you place the first digit in the quotient?

An estimation, like 600 divided by 6, will show you that the answer should be around 100, not 10 or 1000.

Why do so many people gamble and the probability of winning is very low?

People often look at the expected value of the outcome rather than only the winning probability. So with bigger prizes, the probability of winning can be lower and still remain attractive.

Many people experience a buzz out of gambling and it can be very addictive.

Also, there are some games where the probability of winning can be increased - legally. A good poker player, for example can expect to win at a table full of less expert players.

Why does the standard deviation weight extreme deviations more than does the arithmetic mean?

Because the standard deviation is based on the square root of the sum of the squares of the deviations, and, as a result, the sum of the squares of the deviations puts more weight in outliers than does a simple arithmetic mean.

Note: I wrote this and then had second thoughts, but I'm keeping it in so that someone with more knowledge can weigh in (pun intended). I'm not certain how the arithmetic mean factors into the question. I think the questioner, and definitely this answerer, is confused.

Is probability and p value same?

No. p-values are probabilities but they are not the only ones.

What is the probability of flipping 6 fair coins 1 time and getting all heads?

0.015625

For each coin, the probability of it coming up heads is 0.5. For six coins you multiply 0.5 by itself six times, which is 0.5 to the power of 6. This comes out to 0.015625, or about 0.016.

If you work with fractions, then (1/2)^6 = 1/64, which comes out to the same decimal.

What is the probability of rolling a number greater than 27?

The answer depends on how many dice are rolled, whether or not they are fai, how the numbers thrown are combined. For example, in backgammon, a double allows you to move four times the value shown.

The answer depends on how many dice are rolled, whether or not they are fai, how the numbers thrown are combined. For example, in backgammon, a double allows you to move four times the value shown.

The answer depends on how many dice are rolled, whether or not they are fai, how the numbers thrown are combined. For example, in backgammon, a double allows you to move four times the value shown.

The answer depends on how many dice are rolled, whether or not they are fai, how the numbers thrown are combined. For example, in backgammon, a double allows you to move four times the value shown.

How many different ways can the letter WRITTEN be arranged?

7!/2! = 7*6*5*4*3 = 2520

7!/2! = 7*6*5*4*3 = 2520

7!/2! = 7*6*5*4*3 = 2520

7!/2! = 7*6*5*4*3 = 2520

What is the probability of drawing a card from each of 4 decks and getting the nine of clubs from all decks?

the probablility of drawing a nine or clubs from all four randomly shuffled decks with 52 cards is 1 out of 7,311,616

Explain how to find the absolute value of an integer?

The absolute value of an integer is found by ignoring its sign (plus or minus).

Probability of rolling a pair of dice and getting a total of 12?

to get a 12 you need to roll two a six on each of the two dices. the probability of rolling a six 1/6 on each, and to get six on two, is the 1/6 *1/6 = 1/36

What is the probability of drawing 3 red cards with the same value one after the other without replacement given a deck with 13 red cards 13 blue cards 13 black and 13 green all numbered 1 to13?

As described, the deck contains 52 cards, numbered 1 to 13, four times in four colors, red, blue, black, and green. The probability of drawing more than one red card with the same number on it is zero.

What is the probability of selecting a ten or a king of fifty two card decks?

A standard 52 cards deck contains 4 kings and 4 tens. Given that the type of the card does not matter, we have a total of 8 valid cards (4 kings + 4 tens) to choose from a 52 cards deck. Hence the probability is 8/52.

What is probability of drawing a card other than ace?

The probability of drawing a card other than an Ace, in a standard deck of 52 cards, is 48 in 52, or 12 in 13, or about 0.923.

Remember: Probability is simply the number of expected events divided by the number of possible events. In this case, you can count the other events (48 in 52) or you can subtract the results of the negative events (4 in 52) from 1, i.e. 1-(4/52) or 1-(1/13) is still about 0.923.

Applying to suny Albany you hear this school is getting tougher to get into as a freshmen. My stats are 90 GPA top 25 of my class and a 1030 sat which worries me some what do think of my chances?

I got accepted Early Action and I have a 90 GPA, 1070 on SAT's but my Act score was a 31. Yeah just talk to one of your teachers and see if they are an alumni and donate to the college and have them give you a Rec letter, that way you are almost guaranteed a spot

Trending Questions
What is the probability of rolling a four on a die? A red card is drawn from a bridge deck of 52 cards 13 cards are drawn and found to be the same color how many ways you will obtain all black and then the number of way they will be all red? What does raising the standard mean? A bag containing 4 green 6 black 7 white balls a ball is drawn at randomly probability of finding 2 red balls? Can a negative z score be greater than a positive z score? Jobs that use probability? When Jane throws a piece of paper in the bin the probability the piece of paper lands in the bin is 0.6 what is the probability jane misses the bin? How does increasing the sample size affect the sample error of the mean? Can two events be both mutually exclusive and independent at the same time? In a study using 9 samples and in which the population variance is unknown the distribution that should be used to calculate confidence intervals is? What is the probability that two heterozygous parents would have an offspring that produced round seeds from Mendel? What is the probability of rolling a 7 twice in two rolls of a pair of dice? What is the relationship between the probability of an event's occurrence and the probability that it will not occur? What is the probability that the dealer will not get blackjack? What is the probability of drawing a club three through six on the first draw replacing it and drawing a ten card on the second draw? When do you add a probability? What is the probability of getting 100 percent on a four question quiz? Why is the probability of rolling a sum of 2 the same as the probability of rolling a sum of 12? What is the probability of drawing 1 green then 1 blue marble from a bag with 5 blue marbles 5 red marbles 3 green marbles 2 white marbles? How do you Find the mean with a random variable and a probability?